[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jul 24, 2017

Christopher Mello - NOAA Federal christopher.mello at noaa.gov
Mon Jul 24 13:11:03 UTC 2017


Monica,

It looks like outage was earlier than they advertised.  I am not sure if we
will get another outage at 1615 UTC.


Chris





Topic: GOES-16 ABI product outage and mode change scheduled for July 24,
2017

Date/Time Issued: July 21, 2017  2000 UTC

Product(s) or Data Impacted: GOES-16 ABI outage and mode change

Date/Time of Initial Impact: July 24, 2017 1615 UTC and 1900 UTC,
respectively

Date/Time of Expected End: July 24, 2017 1630 UTC and 1915 UTC, respectively

Length of Outage: Approximate 5 minute outage followed by 10 minute
Continuous Full Disk mode

Details/Specifics of Change:

Two ABI calibration scans (desert monitoring) will occur from 1615-1630 UTC
and 1900-1915 UTC on July 24, 2017 following the schedule below:



ABI calibration scan: 1615-1630 UTC - no ABI products generated or
distributed

ABI Continuous Full Disk Mode (mode 4): 1615-1620 UTC

ABI Flex Mode (mode 3) resumes: 1630 UTC



ABI calibration scan: 1900-1905 UTC - no ABI products generated or
distributed

ABI Continuous Full Disk Mode (mode 4): 1900-1915 UTC

ABI Flex Mode (mode 3) resumes: 1915 UTC



Contact Information for Further Information: ESPC Operations at
ESPCOperations at noaa.gov or 301-817-3880 <(301)%20817-3880>





On Mon, Jul 24, 2017 at 9:00 AM, Monica Bozeman - NOAA Federal <
monica.bozeman at noaa.gov> wrote:

> Hi All,
>
> I had sent out an email last week about the GOES-16 outage. This was a
> planned outage from NESDIS for today starting around 10Z and no end time
> was identified. I will forward an update email from NESDIS when I receive
> one.
>
> Monica
>
>
> Monica Bozeman
> Meteorologist/Programmer
> National Hurricane Center
> 301-683-1553 <(301)%20683-1553>
> monica.bozeman at noaa.gov
> <http://www.wdtb.noaa.gov/badges/monica.bozeman>
>
> On Mon, Jul 24, 2017 at 8:54 AM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
>
>> Here is today's summary of NHC computer operations:
>> --- Pablo
>>
>> ----------
>>
>> *TAFB:*
>>
>> 1.) NWPS was started nearly simultaneously for EPAC and ATL on Sunday
>> causing the workstation to crash. Jeff sent email to staff explaining
>> proper procedure to prevent this. TAFB will run this morning one at a time
>> and Doug will monitor to make sure it goes through fine.
>> 2.) Ballyhoo system clock about 100s slower than lx2. NCO Ticket opened
>> on this.
>> 3.) Issue with GFE Sport Forecast. TAFB sent email to Mello who will
>> investigate.
>> 4.) A bug was discovered with TAFB TCMWindTool. Jeff L. fixed it.
>> 5.) 7/21 15Z - 0820L - Performed a satellite fix on 90E invest using
>> GOES-E and a satellite image was not created in conjunction with doing the
>> fix. Matt Sardi to investigate.
>> 1420L - Did not receive a picture for either Greg or 99E after doing the
>> 18Z fixes. Monica and/or Craig to look into this.
>> *HSU:*
>>
>> 1.) GOES-16 missing since around 10Z or so in both AWIPS and NAWIPS. It
>> is coming on the WFO side. Mello and Doug Gaer looking into it. NCO
>> Immediate Level 1 Priority opened with NCO.
>> 1.) Was not able to get FSSE tracker with 24/00Z cycle. Checked this
>> morning and Robbie could not see it either. Monica and/or Craig will look
>> into it.
>> 2.) Mirror screen of AWIPS 2 workstation (lxd-nhcn) pixelated. Chris
>> Mellow fixed it this morning.
>> 3.) Lot of missing NRL MW data. Monica will look into it.
>> 4.) The AMSU size estimates in the SWHG directory are missing. (nothing
>> in the fields last 24 hours) as of 7/21 - 16Z. Monica will look into it.
>> *Other items:*
>>
>> *NHC RFCs:*
>>
>> 1.) RFC for NetApp Migration will be done this week.
>>
>>
>


-- 
Chris Mello
National Hurricane Center
TSB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170724/a76447df/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list